OpenStudy (anonymous):

a college student earned $7800 during summer vacation working as a waiter in a popular restaurant.The student invested part of money at 10% and the rest at 9%.If the student received a total of $740 in interest at the end of the year,how much was invested at 10%

OpenStudy (amistre64):

ax +by = z; a = t-b tx -bx +by = z b(y-x) = z-tx b@ y% = (z-tx)/(%-x) amount@10% = (740/7800 - 7800(.09))/(.10 - .09)

OpenStudy (amistre64):

i prolly missed that in teh z part ...

OpenStudy (anonymous):

not 4400 its 4200

OpenStudy (amistre64):

amount@10% = (740 - 7800(.09))/(.10 - .09) maybe

OpenStudy (amistre64):

3800@10% dbl chk

OpenStudy (anonymous):

i have 3800

OpenStudy (amistre64):

3800(.10) + 4000(.09) ?= 740 380 + 360 ?= 740 740 = 740

OpenStudy (anonymous):

yeah ur right amistre....i did calculation wrong !

OpenStudy (anonymous):

ok